for ba(oh)2 eqwt=1l2 gfwt if we put N instade of eqwt and F instade of gfwt get N=Fl2 but the rule is F=Nl2 describe this

for ba(oh)2
eqwt=1l2 gfwt
if we put N instade of eqwt and F instade of gfwt get
N=Fl2 but the rule is F=Nl2 describe this

Grade:6

2 Answers

Sunil Kumar FP
10 years ago
actually normality and formality are inversly proportional to their equivalent and molecular wt.respectively
For Ba(OH)2 suppose weight taken=w gram volume of the solution=v lt
notmality,N=w/eq.wt*v and formality,F=w/mwt*v
SINCE FOR Ba(OH)2,eqwt=fwt/2
therefore,N=2*F
Sunil Kumar FP
10 years ago
actually normality and formality are inversly proportional to their equivalent and molecular wt.respectively
For Ba(OH)2 suppose weight taken=w gram volume of the solution=v lt
notmality,N=w/eq.wt*v and formality,F=w/mwt*v
SINCE FOR Ba(OH)2,eqwt=fwt/2
therefore,N=2*F
sunil kumar
ask iitian